Complete the magic square
6 ___ ___
27 15 ___
___ ___ 24

Answer by richwmiller(17219)

6 21 18
27 15 3
12 9 24
6+15+24=45
27+15=42
45-42=3
start with the line that has three numbers
then find a line that has two
etc